OpenStudy (anonymous):

find the vertical asymptotes f(x) =x(x-1)/ x^3+16x

myininaya (myininaya):

so you have \[f(x)=\frac{x(x-1)}{x^3}+16x?\]

myininaya (myininaya):

or did you mean to say f(x)=x(x-1)/(x^3+16x)?

myininaya (myininaya):

\[f(x)=\frac{x(x-1)}{x^3+16x}\]?

OpenStudy (anonymous):

16x goes with x^3 on the bottom

myininaya (myininaya):

so that last one I wrote?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok since the vertical asymptote is somewhere the function is not defined then let's look for when the bottom is 0 (we need to be careful with this because we don't want to include a hole as a vertical asymptote) Anyway first step is to factor bottom

OpenStudy (anonymous):

x(x^2+16)